• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name brads

    Log entry time 21:01:23 on March 12, 2012

    Entry number 366034

    keyword=Run plan for Tuesday, Wednesday, Thursday

    Watch for follow-ups to this entry. They may update or supersede what is written below.

    Run Plan for Monday SWING -> Tuesday DAY

    1. Wait until the TE is complete.
    2. HARP scan to determine if amplifier replacement for 1HA05 restores signal.
    3. (30--45min) Sieve data runs with both arms at the production setting (2.228 GeV/c). Do NOT change any spectrometer or other magnet field settings. Take 30--45min of data on carbon foil for both arms, you will need to adjust the prescales to get the CODA rates back up. The target nose will need to be drained of He before doing this measurement.
    4. (10min) Bleedthrough measurement using HRS singles as our metric. Doug will followup with a specific procedure.
    5. (2 hours) BCM calibration (Alex). 0 -- 200nA in 25nA steps -- let's get this done and analyzed ASAP, and update the coefficients everywhere so we can remove some of the beam current ambiguity that we've been dealing with.
    6. (4 hours) Moller measurement (Sasha will run this, Yves will support tune in MCC if needed).
    7. (30 min) Escorted Controlled Access to remove sieves from both arms.
    8. BPM calibration (Alex)
    9. BCM - Tungsten calorimeter calibration run (Alex)
    10. Do a bleedthrough check.
    11. Inelastic data when the above is done until Tuesday DAY. Runplan Here.
    12. Do a bleedthrough check just before 0700.
    13. At 0700 Request a survey to go to Restricted Access.

    General Notes

    Tuesday DAY -> Wednesday OWL

    1. MCC will spend Day shift running Beam studies. We hope to receive beam again in early SWING.
    2. Hall A will be open during DAY shift to move the chicane to the Straight Through position. The target orientation will not be changed.
    3. We will move the Hall to Controlled Access as soon as the chicane move is complete so the the target group can run a TE measurement. IF you wish to work in the Hall, you must be authorized by the Shift Leader.
    4. When beam returns (SWING shift) we will begin with the Optics study (Jixie).
    5. Remaining beam will be dedicated to taking dilution and target background data.
    6. At 0700 Request a survey to go to Restricted Access.

    Wednesday DAY

    1. Hall A will be open during DAY shift to move the chicane to Position #5 (2.2 GeV, Target field = 2.5 T, Target orientation = 90 deg).
    2. TE ?
    3. When beam returns (SWING shift) we will begin g2p production!

    A copy of this log entry has been emailed to: jpchen,slifer,camsonne,doug,rom,brads,jixie,rom,glamazdi,melissac,gilman,rbziel,